OpenStudy (anonymous):

A bag contains 2 white marbles and 7 purple marbles. Two marbles are drawn at random. One marble is drawn and not replaced. Then a second marble is drawn. a. What is the probability of selecting a purple marble and then a white marble? b. What is the probability of selecting two white marbles? c. Is there a greater chance of selecting two white marbles in a row or two purple marbles in a row? Show your work.

OpenStudy (anonymous):

how many marbles total?

OpenStudy (anonymous):

9

OpenStudy (anonymous):

how many purple?

OpenStudy (anonymous):

7

OpenStudy (anonymous):

what is the probability of selecting a purple one first?

OpenStudy (anonymous):

7/9?

OpenStudy (anonymous):

yes, good. now hold on to that number, we are going to need it again

OpenStudy (anonymous):

now that you have selected a purple marble, how many marbles are left?

OpenStudy (anonymous):

6 purple marbles left, and 8 total marbles left.

OpenStudy (anonymous):

right but we need to know how many WHITE marbles are left, because now we have to compute the probability that the next marble selected is white

OpenStudy (anonymous):

2 white marbles are left

OpenStudy (anonymous):

good, so what is the probability that the next marble selected is white?

OpenStudy (anonymous):

2/8?

OpenStudy (anonymous):

exactly (better known as \(\frac{1}{4}\) )

OpenStudy (anonymous):

to finish the problem, multiply the two probabilities i.e. compute \[\frac{7}{9}\times \frac{1}{4}\]

OpenStudy (anonymous):

so that's 7/36?

OpenStudy (anonymous):

it is the probability that the first marble chosen is purple \(\frac{7}{9}\) times the probability that the second marble chosen in white GIVEN that the first one was purple \(\frac{1}{4}\) yes , \(\frac{7}{36}\) is right

OpenStudy (anonymous):

procedure for the probability that both are white is the same, although the answer is of course different

OpenStudy (anonymous):

let me know if you need help with that one too the idea is identical

OpenStudy (anonymous):

I think I have the answer for this part: b. What is the probability of selecting two white marbles? Is it this: 2/9*1/8=1/36, right?

OpenStudy (anonymous):

@satellite73

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

Ok I have part c(the last part finally) Can you help please, then I promise I will leave you alone, you are probably getting tired of me: c. Is there a greater chance of selecting two white marbles in a row or two purple marbles in a row? Show your work. @satellite73

OpenStudy (anonymous):

no not tired of you

OpenStudy (anonymous):

first off it is entirely obvious that it is more likely to get two purple than two white, because there are more purple

OpenStudy (anonymous):

That is what i was thinking

OpenStudy (anonymous):

I just wanted to make sure

OpenStudy (anonymous):

but we know now how to compute both probabilities two white you just computed and got \(\frac{2}{9}\times \frac{1}{8}=\frac{1}{36}\)

OpenStudy (anonymous):

Yep

OpenStudy (anonymous):

two purple is \[\frac{7}{9}\times \frac{6}{8}=\frac{7}{12}\]

OpenStudy (anonymous):

So, two white is: 2/9 x 1/8 = 1/36

OpenStudy (anonymous):

right

OpenStudy (anonymous):

Ok, I got it, thanks! @satellite73

OpenStudy (anonymous):

yw